# Hilbert-Guided Block-Sparse Local Attention

### v1

@@ Abstract
The quadratic compute and memory costs of global self-attention severely limit its use in high-resolution images. Local attention reduces complexity by restricting attention to neighborhoods. Block-sparse kernels can further improve the efficiency of local attention, but conventional local attention patterns often fail to deliver significant speedups because tokens within a window are not contiguous in the 1D sequence. This work proposes a novel method for constructing windows and neighborhoods based on the Hilbert curve. Image tokens are first reordered along a Hilbert curve, and windows and neighborhoods are then formed on the reordered 1D sequence. From a block-sparse perspective, this strategy significantly increases block sparsity and can be combined with existing block-sparse kernels to improve the efficiency of 2D local attention. Experiments show that the proposed Hilbert Window Attention and Hilbert Slide Attention can accelerate window attention and slide attention by about $4\times$ and $18\times$, respectively. To assess practicality, the strategy is instantiated as the Hilbert Window Transformer and the Hilbert Neighborhood Transformer, both of which achieve end-to-end speedups with minimal accuracy loss. Overall, combining Hilbert-guided local attention with block-sparse kernels offers a general and practical approach to enhancing the efficiency of 2D local attention for images. The code is available at https://github.com/Yunge6666/Hilbert-Local-Attention.
@@ 1 Introduction
In recent years, models based on self-attention mechanisms, especially the Transformer architecture, have achieved significant success in computer vision. However, the computational and memory requirements of global self-attention increase quadratically with sequence length, which severely limits its use in processing high-resolution images vaswani2017attention. To address this problem, local attention restricts each token’s receptive field to its neighborhood and thereby reduces complexity.
Some typical works, such as Swin Transformer liu2021swin and Neighborhood Attention Transformer (NAT) hassani2023neighborhood has shown that local attention can maintain the expressiveness of the model while considerably improving computational efficiency, making it a core focus in current research on efficient models. However, existing local attention methods still focus primarily on algorithm and structure design, while optimization at the kernel level remains limited. This is especially true for widely used 2D attention patterns in vision, such as window, sliding window, and neighborhood attention. Current kernel optimization techniques, for example, FlashAttention dao2022flashattention, are often designed for 1D sequences (e.g., text) or regularly structured sparse patterns and are not well adapted to 2D image local attention.
FlexAttention dong2024flex introduces a more flexible approach to optimizing sparse attention, including local attention on 2D images. As a block-sparse attention framework, FlexAttention breaks down the attention operation into different types of blocks: full blocks, partial blocks, and empty blocks. Empty blocks are skipped in computation, which enables the efficient use of sparsity. Analysis of empty block ratio in FlexAttention (Figure 1) reveals that, for fixed sequence lengths such as 1024, 2048, and 4096, the speedup of attention computation is positively correlated with block sparsity, specifically the percentage of empty blocks. Therefore, a higher ratio of empty blocks reduces both computational and memory costs, leading to higher efficiency. This observation suggests that increasing the proportion of empty blocks is an effective way to accelerate attention computation.
@@
Motivated by these observations, this work proposes a novel method, Hilbert-guided construction of windows and neighborhoods, to enable more efficient use of sparsity by increasing the empty blocks ratio. Specifically, the Hilbert curve hilbert1935stetige is adopted due to its strong locality-preserving property jagadish1997analysis: when mapping 2D image tokens to a 1D sequence, it maintains neighborhood relations. This reordering greatly increases the ratio of empty blocks in the local attention pattern and reduces the ratio of partial blocks. Within block-sparse kernel execution, empty blocks are skipped, directly reducing computational and memory overhead. Moreover, the programmable interface of the FlexAttention framework enables deployment without modifying the model or training pipeline and without hand-written kernels. In short, although local attention is inherently sparse, regular square windows in row-major order still produce many partial blocks that require computation. By remapping 2D local attention patterns into contiguous 1D blocks via the Hilbert reordering, the empty block ratio is greatly increased, leading to a significant acceleration of local attention.
@@ 2.1 Local Attention for Images
In Vision Transformer dosovitskiy2020image, while globally computed full attention mechanisms are effective, they come with high computational costs. For high-dimensional data, such as images, their quadratic complexity becomes a significant bottleneck. Local attention mechanisms are based on a reasonable inductive bias: a visual token typically only needs to interact with tokens in its surrounding neighborhood. This assumption significantly reduces computational complexity, making it a mainstream approach for building efficient vision transformer  vaswani2021halonet,  yang2021focaldong2022cswintu2022maxvit,chen2021regionvit,liu2022swinli2022simvit.
Stand-Alone Self-Attention (SASA) ramachandran2019stand was an early pioneering attempt to entirely replace convolutional layers with local self-attention layers in visual tasks. Swin Transformer liu2021swin introduced a local attention mechanism based on regular windows and a shifted window strategy, cleverly enabling cross-window information exchange. It has become a milestone work in the field of vision transformers. The Slide Transformer pan2023slide revisits sliding window attention by replacing the inefficient $Im2Col$ function with depthwise convolution and equipping it with a learned shift module, thereby achieving efficient local attention. NAT proposed neighborhood attention, which differs from previous sliding window attention in its boundary handling: instead of padding zeros, NAT repeats the same “window” at the boundaries.
@@ 2.2 Attention Kernel Optimization
The goal of efficient attention computation techniques is to directly address the memory and computational bottlenecks of attention mechanisms at the system level. In recent years, breakthroughs in this field have primarily stemmed from the design of hardware-aware kernels zhang2024sageattentionkwon2023efficientliu2023ringyuan2025nativexu2025xattention.
FlashAttention dao2022flashattention is a foundational work. It was the first to restructure attention computation from the perspective of memory access cost, proposing an I/O-aware algorithm. Its core innovation lies in a tiling strategy that decomposes the computation to be performed in the GPU’s high-speed SRAM, avoiding the materialization of large intermediate attention matrices in memory. This reduces the memory complexity and achieves a several-fold speedup. FlashAttentionv2 dao2023flashattention further exploits hardware potential by restructuring parallelization strategies and refining task scheduling, significantly reducing kernel synchronization overhead and achieving higher throughput. The latest FlashAttentionv3 shah2024flashattention begins to leverage features of newer-generation GPUs (e.g., NVIDIA H100) to further improve computational efficiency NATTEN also specifically conducted extreme kernel-level optimizations for Neighborhood Attention hassani2024fasterhassani2025generalized. They eliminated the inherent $O(N^{2})$ intermediate memory cost in neighborhood attention, providing a highly efficient implementation for this specific local attention pattern. XFormers xFormers2022 provides a library integrating multiple efficient attention implementations. XFormers includes not only memory-efficient attention implementations but also various predefined patterns.
However, the high-performance kernels of FlashAttention, NATTEN, and XFormers are all hand-optimized for predefined, fixed sparsity patterns. This introduces a key limitation: when research requires experimenting with a completely new, unimplemented attention pattern, significant effort is still needed to redesign and reimplement the underlying kernel, creating a high barrier to entry and a lack of flexibility. To address the flexibility issue that FlexAttention dong2024flex was proposed to solve, a shift in programming paradigm is represented. FlexAttention builds a general compilation framework that allows developers to define arbitrary block-sparse attention patterns directly using high-level Python code. The compiler then automatically generates highly optimized GPU kernel code at compile time. This approach decouples algorithm innovation from low-level optimization. This work leverages the generality of FlexAttention to explore and further optimize efficiency bottlenecks of local attention in vision.
@@ 3.1 Hilbert Local Attention
The proposed Hilbert-guided local attention is illustrated in Figure 3. Given $N$ tokens, attention is viewed as computation on an N×N matrix: rows correspond to each query ($q$) token, and columns correspond to all key ($k$) tokens. The attention pattern of $16$ tokens is visualized as the $16\times 16$ matrix.
@@
On GPUs, the overall runtime of block-sparse attention is jointly determined by the total workload and the effective parallel throughput (nvidia_roofline_blog). It can be approximated as follows:
@@
where $\alpha$ represents the overhead per CTA (compute thread array), covering kernel/CTA launch, $q$ block loading, and initialization (cuda_runtime_occupancy_api); $\beta$ denotes the unit cost of processing a single non-empty block, which includes loading $q$ block and $k/v$ block, computing $qk^{\top}$ within the block, score modification (e.g., relative position bias or Alibi press2021train), online softmax dao2022flashattention, and aggregation with value ($v$); $r_{i}$ is the number of non-empty blocks contained in the $i_{th}$ CTA; and $P_{\text{eff}}$ is the effective parallelism, determined by the number of streaming multiprocessors (SMs) and some memory factors (nsight_compute_profiling_guide). By skipping empty blocks, block-sparse attention avoids launching corresponding CTA and loading key/value data, thereby reducing computation and memory access. The proposed Hilbert reordering further accelerates block-sparse attention by increasing the ratio of empty blocks.
@@ 3.2 Hilbert Window Transformer
WSA is applied in Swin Transformer, and it does not leverage block sparsity; instead, Swin Transformer performs dense WSA within each window. Although this approach implicitly partitions the sequence into windows, it does not further optimize the underlying computation. Block-sparse kernel can accelerate WSA; however, it may sometimes suffer from efficiency loss due to generating more partial blocks. By applying HWA, the number of empty blocks increases and the number of partial blocks decreases, thereby better leveraging the potential of block-sparse attention. Based on this idea, this work introduces the Hilbert Window Transformer (HWT). HWT shares a similar architecture with Swin Transformer, with the main differences lying in the window construction, the shift window operation, and the use of RPB. As shown in Figure 4, the image is first divided into tokens (or patches), which are then reordered according to the Hilbert curve path. Since feature maps of the same size generate the same Hilbert curve path, the path can be precomputed and cached for reuse in subsequent feature maps of identical dimensions.
@@ 4.1 Implementation Details
Model throughput and runtime depend on the hardware platform and software stack. Different GPU models, CUDA versions, or PyTorch versions can result in varying speedups. All experiments here are conducted on RTX 3080 GPU using CUDA 12.6 and PyTorch 2.7.0. For reference, the Appendix also reports results on an A100. The goal is to verify that Hilbert reordering increases the empty block ratio and thereby improves the efficiency of image local attention. Because the empty block ratio is mainly affected by input size, window/kernel size, and block size, we evaluate across different inputs, window sizes, and block sizes, comparing HWA with WSA and HSA/HNA with SA and NA. In addition, the implementation of WSA/SA/NA in FlexAttention is also evaluated and compared. We further include evaluations that combine input processing steps, such as window partitioning, Hilbert reordering, and QKV projection, with the attention computation. Finally, to assess feasibility in full models, we report the accuracy and efficiency of HWT and HNT on ImageNet deng2009imagenet. Due to variability in CUDA timing, all evaluations are averages over 10 runs, each run consisting of 100 iterations with $25\%$ warm-up and $75\%$ measurement.
@@ 4.2 Results and Analysis
Window Attention Result. Table 1 reports the efficiency of WSA, WSA (Flex), and HWA under different input feature-map sizes and window sizes. WSA uses a dense kernel as the baseline. WSA (Flex) calls the block-sparse kernel in FlexAttention, but because it contains many partial blocks that require element-wise masking, it runs slower than WSA in all settings. The sparsity metric denotes the ratio of empty blocks, and HWA shows higher sparsity than WSA (Flex) in every configuration. As a result, HWA is faster than WSA (Flex) for both inference and training, and achieves up to $4.0\times$ speedup over the dense WSA. Benefiting from FlexAttention’s integration of FlashAttention, both WSA (Flex) and HWA use much less memory than WSA in forward and backward. Memory consumption is dominated by the size of the QKV tensors and is only weakly affected by block size and window size. Thus, for the same input, WSA (Flex) and HWA have the same memory usage. For example, when the input size is I=96, their memory stays unchanged across different window sizes. In contrast, WSA’s memory grows with input and window size because dense computation materializes intermediate attention results within each window, whereas FlexAttention stores only lightweight block mask metadata and does not keep the full attention matrix. Overall, WSA (Flex) and HWA reduce training and inference memory compared with WSA, and HWA further improves computation speed.
@@
Attention
Block
Size
Forward Sparsity
@@
Before computing attention, the input is processed with two steps: a reshape and linear projections to Q, K, and V. In WSA, the reshape is window partitioning. In HWA, the reshape is Hilbert reordering. WSA (Flex) does not require a reshape because FlexAttention operates directly on QKV obtained from the linear projection of the features. Figure 6(a) reports results of window attention variants for input $128\times 128$ with a window $16\times 16$. The attention time here differs from Table 1 because this evaluation includes RPB, but HWA remains faster than WSA. In the QKV projection stage, HWA is also faster than WSA. WSA first partitions windows, producing many small matrices and incurring more kernel launches and memory traffic, whereas HWA processes fewer, larger matrices, which improves throughput. WSA (Flex) and HWA both project the full input feature map, so their QKV projection times are essentially the same. Overall, HWA takes 1.63 ms, which is $2.6\times$ faster than WSA’s 4.22 ms and $3.9\times$ faster than WSA (Flex).
@@
Slide/Neighborhood Attention Results. Table 3 reports results for SA and NA under different input sizes and kernel sizes. SA is an unoptimized baseline, and its runtime and memory usage are much higher than those of attention variants with an optimized kernel. NA2D and HNA can be implemented with either FlexAttention or NATTEN. Under the same settings, HSA shows higher sparsity and larger speedups than SA (Flex), and HNA (Flex) shows higher sparsity and better speed than NA2D (Flex). With the NATTEN kernel, HNA (NAT) is also faster than NA2D. Concretely, at input $56\times 56$ with kernel size $7\times 7$, HNA (NAT) is $48.8\times$ faster than SA and $1.8\times$ faster than NA2D. At input $96\times 96$ with kernel size $7\times 7$, SA runs out of memory (OOM), and HSA and HNA are about $1.8–2.1\times$ faster than NA2D (NAT). Figure 6(b) presents an evaluation that combines attention with input processing. HSA and HNA require a reshape step, specifically Hilbert reordering, whereas the other methods project the input directly to QKV and then compute attention. SA’s neighborhood construction is both memory-intensive and time-consuming, making it non-competitive. Under the $I=56\times 56$ and $K=7\times 7$, NA2D (NAT) and HNA (NAT) have similar runtimes. At higher resolutions and larger kernel sizes, HNA and HSA are faster than NA2D.
@@
HWT and HNT Results. Training was conducted on 8 V100 GPUs for HWT-T and HNT-mini, with training settings and other hyperparameters matched to Swin-Tiny and NAT-mini. Because V100 does not support FlexAttention well, HWT used a dense kernel during training. Additional correctness tests verified that HWA with the dense kernel produces the same outputs as HWA with FlexAttention. Specifically, the correctness tests ensured that mask_modmask\_mod and score_modscore\_mod were designed so that the same attention pattern has identical outputs for the FlexAttention block-sparse kernel and for the dense implementation. Similar correctness checks were performed for HSA vs. SA and for HNA vs. NAT. Table 4 reports ImageNet-1K accuracy. At image resolutions of $224\times 224$ and $256\times 256$, applying the proposed HWA and HNA in HWT and HNT results in at most a $0.2\%$ drop in accuracy, whereas Figure 7 shows that both models deliver substantial efficiency gains, especially at high resolution. These results confirm the feasibility of the proposed HWA and HNA.

### v2

# Hilbert-Guided Sparse Local Attention

@@ Abstract
The quadratic compute and memory costs of global self-attention severely limit its use in high-resolution images. Local attention reduces complexity by restricting attention to neighborhoods. Block-sparse kernels can further improve the efficiency of local attention, but conventional local attention patterns often fail to deliver significant speedups because tokens within a window are not contiguous in the 1D sequence. This work proposes a novel method for constructing windows and neighborhoods based on the Hilbert curve. Image tokens are first reordered along a Hilbert curve, and windows and neighborhoods are then formed on the reordered 1D sequence. From a block-sparse perspective, this strategy significantly increases block sparsity and can be combined with existing block-sparse kernels to improve the efficiency of 2D local attention. Experiments show that the proposed Hilbert Window Attention and Hilbert Slide Attention can accelerate window attention and slide attention by about $4\times$ and $18\times$, respectively. To assess practicality, the strategy is instantiated as the Hilbert Window Transformer and the Hilbert Neighborhood Transformer, both of which achieve end-to-end speedups with minimal accuracy loss. Overall, combining Hilbert-guided local attention with block-sparse kernels offers a general and practical approach to enhancing the efficiency of 2D local attention for images.
footnotetext: The code is available at: https://github.com/Yunge6666/Hilbert-Local-Attention
@@ 1 Introduction
In recent years, models based on self-attention mechanisms, especially the Transformer architecture, have achieved significant success in computer vision. However, the computational and memory requirements of global self-attention increase quadratically with sequence length, which severely limits its use in processing high-resolution images (Vaswani et al., 2017). To address this problem, local attention restricts each token’s receptive field to its neighborhood and thereby reduces complexity.
Some typical works, such as Swin Transformer (Liu et al., 2021) and Neighborhood Attention Transformer (NAT) (Hassani et al., 2023) has shown that local attention can maintain the expressiveness of the model while considerably improving computational efficiency, making it a core focus in current research on efficient models. However, existing local attention methods still focus primarily on algorithm and structure design, while optimization at the kernel level remains limited. This is especially true for widely used 2D attention patterns in vision, such as window, sliding window, and neighborhood attention. Current kernel optimization techniques, for example, FlashAttention (Dao et al., 2022), are often designed for 1D sequences (e.g., text) or regularly structured sparse patterns and are not well adapted to 2D image local attention.
FlexAttention (Dong et al., 2024) introduces a more flexible approach to optimizing sparse attention, including local attention on 2D images. As a block-sparse attention framework, FlexAttention breaks down the attention operation into different types of blocks: full blocks, partial blocks, and empty blocks. Empty blocks are skipped in computation, which enables the efficient use of sparsity. Analysis of empty block ratio in FlexAttention (Figure 1) reveals that, for fixed sequence lengths such as 1024, 2048, and 4096, the speedup of attention computation is positively correlated with block sparsity, specifically the percentage of empty blocks. Therefore, a higher ratio of empty blocks reduces both computational and memory costs, leading to higher efficiency. This observation suggests that increasing the proportion of empty blocks is an effective way to accelerate attention computation.
@@
Motivated by these observations, this work proposes a novel method, Hilbert-guided construction of windows and neighborhoods, to enable more efficient use of sparsity by increasing the empty blocks ratio. Specifically, the Hilbert curve (Hilbert, 1935) is adopted due to its strong locality-preserving property (Jagadish, 1997): when mapping 2D image tokens to a 1D sequence, it maintains neighborhood relations. This reordering greatly increases the ratio of empty blocks in the local attention pattern and reduces the ratio of partial blocks. Within block-sparse kernel execution, empty blocks are skipped, directly reducing computational and memory overhead. Moreover, the programmable interface of the FlexAttention framework enables deployment without modifying the model or training pipeline and without hand-written kernels. In short, although local attention is inherently sparse, regular square windows in row-major order still produce many partial blocks that require computation. By remapping 2D local attention patterns into contiguous 1D blocks via the Hilbert reordering, the empty block ratio is greatly increased, leading to a significant acceleration of local attention.
@@ 2.1 Local Attention for Images
In Vision Transformer (Dosovitskiy et al., 2020), while globally computed full attention mechanisms are effective, they come with high computational costs. For high-dimensional data, such as images, their quadratic complexity becomes a significant bottleneck. Local attention mechanisms are based on a reasonable inductive bias: a visual token typically only needs to interact with tokens in its surrounding neighborhood. This assumption significantly reduces computational complexity, making it a mainstream approach for building efficient vision transformer  (Vaswani et al., 2021),  (Yang et al., 2021)(Dong et al., 2022)(Tu et al., 2022),(Chen et al., 2021),(Liu et al., 2022)(Li et al., 2022).
Stand-Alone Self-Attention (SASA) (Ramachandran et al., 2019) was an early pioneering attempt to entirely replace convolutional layers with local self-attention layers in visual tasks. Swin Transformer (Liu et al., 2021) introduced a local attention mechanism based on regular windows and a shifted window strategy, cleverly enabling cross-window information exchange. It has become a milestone work in the field of vision transformers. The Slide Transformer (Pan et al., 2023) revisits sliding window attention by replacing the inefficient $Im2Col$ function with depthwise convolution and equipping it with a learned shift module, thereby achieving efficient local attention. NAT proposed neighborhood attention, which differs from previous sliding window attention in its boundary handling: instead of padding zeros, NAT repeats the same “window” at the boundaries.
@@ 2.2 Attention Kernel Optimization
The goal of efficient attention computation techniques is to directly address the memory and computational bottlenecks of attention mechanisms at the system level. In recent years, breakthroughs in this field have primarily stemmed from the design of hardware-aware kernels (Zhang et al., 2024)(Kwon et al., 2023)(Liu et al., 2023)(Yuan et al., 2025)(Xu et al., 2025).
FlashAttention (Dao et al., 2022) is a foundational work. It was the first to restructure attention computation from the perspective of memory access cost, proposing an I/O-aware algorithm. Its core innovation lies in a tiling strategy that decomposes the computation to be performed in the GPU’s high-speed SRAM, avoiding the materialization of large intermediate attention matrices in memory. This reduces the memory complexity and achieves a several-fold speedup. FlashAttentionv2 (Dao, 2023) further exploits hardware potential by restructuring parallelization strategies and refining task scheduling, significantly reducing kernel synchronization overhead and achieving higher throughput. The latest FlashAttentionv3 (Shah et al., 2024) begins to leverage features of newer-generation GPUs (e.g., NVIDIA H100) to further improve computational efficiency NATTEN also specifically conducted extreme kernel-level optimizations for Neighborhood Attention (Hassani et al., 2024)(Hassani et al., 2025). They eliminated the inherent $O(N^{2})$ intermediate memory cost in neighborhood attention, providing a highly efficient implementation for this specific local attention pattern. XFormers (Lefaudeux et al., 2022) provides a library integrating multiple efficient attention implementations. XFormers includes not only memory-efficient attention implementations but also various predefined patterns.
However, the high-performance kernels of FlashAttention, NATTEN, and XFormers are all hand-optimized for predefined, fixed sparsity patterns. This introduces a key limitation: when research requires experimenting with a completely new, unimplemented attention pattern, significant effort is still needed to redesign and reimplement the underlying kernel, creating a high barrier to entry and a lack of flexibility. To address the flexibility issue that FlexAttention (Dong et al., 2024) was proposed to solve, a shift in programming paradigm is represented. FlexAttention builds a general compilation framework that allows developers to define arbitrary block-sparse attention patterns directly using high-level Python code. The compiler then automatically generates highly optimized GPU kernel code at compile time. This approach decouples algorithm innovation from low-level optimization. This work leverages the generality of FlexAttention to explore and further optimize efficiency bottlenecks of local attention in vision.
@@ 3.1 Hilbert Local Attention
The proposed Hilbert-guided local attention is illustrated in Figure 3. Given $N$ tokens, attention is viewed as computation on an N×N matrix: rows correspond to each query ($q$) token, and columns correspond to all key ($k$) tokens. The attention pattern of $16$ tokens is visualized into the $16\times 16$ matrix.
@@
On GPUs, the overall runtime of block-sparse attention is jointly determined by the total workload and the effective parallel throughput (Blog, 2020). It can be approximated as follows:
@@
where $\alpha$ represents the overhead per CTA (compute thread array), covering kernel/CTA launch, $q$ block loading, and initialization (Corporation, 2025a); $\beta$ denotes the unit cost of processing a single non-empty block, which includes loading $q$ block and $k/v$ block, computing $qk^{\top}$ within the block, score modification (e.g., relative position bias or Alibi (Press et al., 2021)), online softmax (Dao et al., 2022), and aggregation with value ($v$); $r_{i}$ is the number of non-empty blocks contained in the $i_{th}$ CTA; and $P_{\text{eff}}$ is the effective parallelism, determined by the number of streaming multiprocessors (SMs) and some memory factors (Corporation, 2025b). By skipping empty blocks, block-sparse attention avoids launching corresponding CTA and loading key/value data, thereby reducing computation and memory access. The proposed Hilbert reordering further accelerates block-sparse attention by increasing the ratio of empty blocks.
@@ 3.2 Hilbert Window Transformer
WSA is applied in Swin Transformer, and it does not leverage block sparsity; instead, Swin Transformer performs dense WSA within each window. Although this approach implicitly partitions the sequence into windows, it does not further optimize the underlying computation. Block-sparse kernel can accelerate WSA; however, it may sometimes suffer from efficiency loss due to generating more partial blocks. By applying HWA, the number of empty blocks increases and the number of partial blocks decreases, thereby better leveraging the potential of block-sparse attention. Based on this idea, this work introduces the Hilbert Window Transformer (HWT). HWT shares a similar architecture with Swin Transformer, with the main differences lying in the window construction, the shift window operation, and the use of RPB. As shown in Figure 4, the image is first divided into tokens (or patches), which are then reordered according to the Hilbert curve path. Since feature maps with the same size produce the same Hilbert-curve path, a feature map of size $(H,W)$ corresponds to a unique path $P_{(H,W)}$. The path $P_{(H,W)}$ is computed and cached at model initialization, and whenever a feature map with the same size $(H,W)$ is processed, the cached path $P_{(H,W)}$ can be reused.
@@ 4.1 Implementation Details
Model throughput and runtime depend on the hardware platform and software stack. Different GPU models, CUDA versions, or PyTorch versions can result in varying speedups. All experiments here are conducted on RTX 3080 GPU using CUDA 12.6 and PyTorch 2.7.0. For reference, the Appendix also reports results on an A100. The goal is to verify that Hilbert reordering increases the empty block ratio and thereby improves the efficiency of image local attention. Because the empty block ratio is mainly affected by input size, window/kernel size, and block size, we evaluate across different inputs, window sizes, and block sizes, comparing HWA with WSA and HSA/HNA with SA and NA. In addition, the implementation of WSA/SA/NA in FlexAttention is also evaluated and compared. We further include evaluations that combine input processing steps, such as window partitioning, Hilbert reordering, and QKV projection, with the attention computation. Finally, to assess feasibility in full models, we report the accuracy and efficiency of HWT and HNT on ImageNet (Deng et al., 2009). To show that HWA/HSA/HNA can be adapted to different optimized kernels without hand-written kernels, we further evaluate and compare them by applying FlashAttentionv2, xFormers, and NAT to them. Due to variability in CUDA timing, all evaluations are averages over 10 runs, each run consisting of 100 iterations with $25\%$ warm-up and $75\%$ measurement.
@@ 4.2 Results and Analysis
Window Attention Result. Table 1 reports the efficiency of WSA and HWA with different optimized kernels under different input feature-map sizes and window sizes. WSA uses a dense kernel as the baseline. WSA (Flex) calls the block-sparse kernel in FlexAttention, but because it contains many partial blocks that require element-wise masking, it runs slower than WSA in all settings. The sparsity metric denotes the ratio of empty blocks, and HWA (Flex) shows higher sparsity than WSA (Flex) in every configuration. As a result, HWA (Flex) is faster than WSA (Flex) for both inference and training, and achieves up to $4.0\times$ speedup over the dense WSA. Benefiting from FlexAttention’s integration of FlashAttentionv2, both WSA (Flex) and HWA (Flex) use much less memory than WSA in forward and backward. Memory consumption is dominated by the size of the QKV tensors and is only weakly affected by block size and window size. Thus, for the same input, WSA (Flex) and HWA (Flex) have the same memory usage. For example, when the input size is I=96, their memory stays unchanged across different window sizes. In contrast, WSA’s memory grows with input and window size because dense computation materializes intermediate attention results within each window, whereas FlexAttention stores only lightweight block mask metadata and does not keep the full attention matrix. Overall, WSA (Flex) and HWA (Flex) reduce training and inference memory compared with WSA, and HWA (Flex) further improves computation speed. Since FlashAttentionv2 does not support arbitrary masks, WSA and HWA cannot be directly accelerated by it without manually modifying the kernel. Although xFormers supports more attention patterns, its implementation of WSA still uses an element-wise sparse path that stores many individual non-zero elements, which leads to high memory usage and is usually slower. In contrast, HWA can use block-diagonal fused dense kernels, so HWA (xFormers) in Table 1 is significantly faster than WSA (xFormers).
@@
Attention Block Size Forward Sparsity
@@
Before computing attention, the input is processed with two steps: a reshape and linear projections to Q, K, and V. In WSA, the reshape is window partitioning. In HWA, the reshape is Hilbert reordering. WSA (Flex) does not require a reshape because FlexAttention operates directly on QKV obtained from the linear projection of the features. Figure 6(a) reports results of window attention variants for input $128\times 128$ with a window $16\times 16$. The attention time here differs from Table 1 because this evaluation includes RPB, but HWA (Flex) remains faster than WSA. In the QKV projection stage, HWA (Flex) is also faster than WSA. WSA first partitions windows, producing many small matrices and incurring more kernel launches and memory traffic, whereas HWA (Flex) processes fewer, larger matrices, which improves throughput. WSA (Flex) and HWA (Flex) both project the full input feature map, so their QKV projection times are essentially the same. Overall, HWA (Flex) takes 1.63 ms, which is $2.6\times$ faster than WSA’s 4.22 ms and $3.9\times$ faster than WSA (Flex).
Slide/Neighborhood Attention Results. Table 3 reports results for SA and NA under different input sizes and kernel sizes. SA is an unoptimized baseline, and its runtime and memory usage are much higher than those of attention variants with an optimized kernel. SA can use the optimized kernels from xFormers and FlexAttention, while HSA can additionally be accelerated by FlashAttentionv2, because FlashAttentionv2 natively supports sliding-window (banded) patterns that match the HSA pattern. NA2D and HNA can be implemented with either FlexAttention or NATTEN. For SA(xFormers), it follows the same element-wise sparse path as WSA (xFormers) and is therefore relatively slow, whereas HSA can directly use the optimized kernel in xFormers that is specifically designed for banded (sliding-window) attention, so it is much faster than SA (xFormers). Similarly, FlashAttentionv2 also provides a specialized optimized kernel for banded patterns, so HSA (FA2) achieves a very significant improvement in efficiency as well. Under the same settings, HSA (Flex) shows higher sparsity and larger speedups than SA (Flex), and HNA (Flex) shows higher sparsity and better speed than NA2D (Flex). With the NATTEN kernel, HNA (NAT) is also faster than NA2D. Concretely, at input $56\times 56$ with kernel size $7\times 7$, HNA (NAT) is $48.8\times$ faster than SA and $1.8\times$ faster than NA2D. At input $96\times 96$ with kernel size $7\times 7$, SA runs out of memory (OOM), and HSA (Flex) and HNA (NAT) are about $1.8–2.1\times$ faster than NA2D (NAT).
@@
Figure 6(b) presents an evaluation that combines attention with input processing. HSA and HNA require a reshape step, specifically Hilbert reordering, whereas the other methods project the input directly to QKV and then compute attention. SA’s neighborhood construction is both memory-intensive and time-consuming, making it non-competitive. Under the $I=56\times 56$ and $K=7\times 7$, NA2D (NAT) and HNA (NAT) have similar runtimes. At higher resolutions and larger kernel sizes, HNA (NAT) and HSA (Flex) are faster than the corresponding NA2D.
@@
Table 5: Accuracy on CIFAR datasets. Due to the low resolution of the CIFAR dataset, all models in the table are configured with only three layers, with depth=[2,6,4], head number=[3,6,12], and head dimention of 96. The patch size is 2, while the window size for SWIN and HWT is set to 4*4, and the kernel size for NAT is 7*7. The training epoch is 300 for all models.
Model Configuration Window CIFAR10 (32x32) CIFAR100 (32x32)
SWIN D=[2,6,4], H=[3,6,12] 4x4 92.3% 76.6%
HWT D=[2,6,4], H=[3,6,12] 16 92.2% 76.4%
NAT D=[2,6,4], H=[3,6,12] 7x7 94.2% 79.9%
HNT D=[2,6,4], H=[3,6,12] 49 94.1% 79.9%
HWT and HNT Results. Training was conducted on 8 V100 GPUs for HWT-T and HNT-mini, with training settings and other hyperparameters matched to Swin-Tiny and NAT-mini. Because V100 does not support FlexAttention well, HWT used a dense kernel during training. Additional correctness tests verified that HWA with the dense kernel produces the same outputs as HWA with FlexAttention. Specifically, the correctness tests ensured that mask_modmask\_mod and score_modscore\_mod were designed so that the same attention pattern has identical outputs for the FlexAttention block-sparse kernel and for the dense implementation. Similar correctness checks were performed for HSA vs. SA and for HNA vs. NAT. Table 4 reports ImageNet-1K accuracy. At image resolutions of $224\times 224$ and $256\times 256$, applying the proposed HWA and HNA in HWT and HNT results in at most a $0.2\%$ drop in accuracy, whereas Figure 7 shows that both models deliver substantial efficiency gains, especially at high resolution. Table 5 shows the performance of HWT and HNT on CIFAR10 and CIFAR100. Compared with Swin and NAT, the accuracy of HWT and HNT is comparable, with performance differences within $0.2\%$. These results confirm the feasibility of the proposed HWA and HNA. Since HWT and HNT retain standard hierarchical backbone structures that are commonly used in object detection, semantic segmentation, and image or video generation pipelines, the same Hilbert local attention modules can, in principle, be reused in these settings as well. A comprehensive investigation into these downstream tasks will be conducted in future work.